# Neural density functional theory of liquid-gas phase coexistence

### v1

@@ Abstract
We use supervised machine learning together with the concepts of classical density functional theory to investigate the effects of interparticle attraction on the pair structure, thermodynamics, bulk liquid-gas coexistence, and associated interfacial phenomena in many-body systems. Local learning of the one-body direct correlation functional is based on Monte Carlo simulations of inhomogeneous systems with randomized thermodynamic conditions, randomized planar shapes of the external potential, and randomized box sizes. Focusing on the prototypical Lennard-Jones system, we test predictions of the resulting neural attractive density functional across a broad spectrum of physical behaviour associated with liquid-gas phase coexistence in bulk and at interfaces. We analyse the bulk radial distribution function $g(r)$ obtained from automatic differentiation and the Ornstein-Zernike route and determine i) the Fisher-Widom line, i.e. the crossover of the asymptotic (large distance) decay of $g(r)$ from monotonic to oscillatory, ii) the (Widom) line of maximal correlation length, iii) the line of maximal isothermal compressibility and iv) the spinodal by calculating the poles of the structure factor in the complex plane. The bulk binodal and the density profile of the free liquid-gas interface are obtained from density functional minimization and the corresponding surface tension from functional line integration. We also show that the neural functional describes accurately the phenomena of drying at a hard wall and of capillary evaporation for a liquid confined in a slit pore. Our neural framework yields results that improve significantly upon standard mean-field treatments of interparticle attraction. Comparison with independent simulation results demonstrates a consistent picture of phase separation even when restricting the training to supercritical states only. We argue that phase coexistence and its associated signatures can be discovered as emerging phenomena via functional mappings and educated extrapolation.
@@ I Introduction
The emergence of two or more distinct thermodynamic stable phases on varying thermodynamic conditions is arguably one of the most striking phenomena in statistical mechanics, whether this occurs in bulk or at interfaces, in pure or in multi-component systems. Unsurprisingly, the recent surge in the use of machine learning techniques in physics [1, 2] has focused on the prototypical (lattice-based) Ising model for developing appropriate techniques and strategies to investigate phase-separating systems. Such work includes the finite-size analysis in neural network classification of critical phenomena [3] and mapping out phase diagrams with generative classifiers [4]. Detecting the approach to a critical point also constitutes a central task in more general dynamical systems, which commonly requires the application of advanced computational methods [5, 6, 7].
@@
As DFT is an exact formulation of the many-body statistical mechanics, the choice of the excess free energy functional is the only approximation that enters a given study. Once it is specified consistent and complete investigation of a wide variety of properties can be made. Computational limitations, leaving aside some intricacies of implementing non-local treatments for hard spheres [13, 14], are minor, certainly for systems which exhibit planar or spherical symmetry. Hence wide parameter sweeps and close monitoring of the effects of small changes in control parameters, especially near phase coexistence, are readily carried out enabling the investigation of subtle phenomena such as phase transitions at substrates and in confinement.
@@
Machine-learning provides a very different perspective for addressing the limitations of analytical approaches (meaning writing down explicit free energy functionals) by incorporating quasiexact simulation reference data for the construction of functional relationships, as was pursued for the classical [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32] and quantum (electronic) worlds [33, 34, 35, 36, 37, 38, 39, 40]. Amongst the different machine-learning approaches [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41] that were put forward, the neural functional theory based on local one-body learning [24, 25, 26, 27, 28, 29] in inhomogeneous training systems has proved to be a versatile and highly accurate tool, both in equilibrium [25, 26, 27, 28] as well as for microscopically-resolved nonequilibrium flow problems [24, 29]. In the latter case the required functional relationships are those of power functional theory [42, 43].
@@
The neural functional approach was investigated in great detail for models with purely repulsive (hard core) potentials. It provides an excellent approximation for Percus’ exact free energy functional for hard rods in one dimension [26] and was shown to constitute a clear improvement [25] over the already highly accurate White-Bear Mk. II version of fundamental-measure theory [54, 14] for hard spheres in three dimensions. Importantly, the neural functional method is not restricted to hard cores; it applies to general interatomic potentials that include attraction. That the neural functional can treat attraction successfully was illustrated in determinations of the structure of the Lennard-Jones fluid at a fixed supercritical temperature [25, 28]; this constitutes an important test case [23].
These recent investigations did not address the fundamental issue of how the presence of a phase transition might be accounted for within the framework of a neural density functional. Here we focus on the liquid-gas transition which is a basic manifestation of the presence of interparticle attraction and seek to assess whether the neural functional can describe i) phase coexistence and the approach to the associated critical point, ii) surface tension and density profiles of the liquid-gas interface, iii) drying and capillary evaporation transitions that occur at subcritical temperatures and iv) how accurately the approach performs for both bulk and interfacial properties. To the best of our knowledge, none of these issues were addressed in previous machine learning investigations. Specifically, we extend the neural methodology [25, 26] by introducing thermal training and investigate whether neural functionals can describe physical phenomena that occur in the three-dimensional Lennard-Jones system in sub- as well as supercritical regions of the phase diagram. The answer is emphatically: yes.
@@
All subsequent results originate from this neural functional method and are described in Sec. III. An account of the emerging bulk pair correlation structure is given in Sec. III.1. Results for the lines in the phase diagram where the isothermal compressibility and the true correlation length are maximal, together with the Fisher-Widom line and the spinodal, are presented Sec. III.2. Our results for the liquid-gas binodal and estimate of the critical point are described in Sec. III.3. The bulk equation of state, liquid-gas density profiles, and the corresponding surface tension are laid out in Sec. III.4. A description of the divergence of the correlation length in the critical region and corresponding Ornstein-Zernike plots are given in Sec. III.5. Results for inhomogeneous fluids, that describe our predictions for drying at a hard wall, for capillary evaporation in a slit pore and for the behaviour of locally resolved density fluctuations, are presented in Sec. III.6. In Sec. III.7 we compare with results obtained from a neural functional trained with data from supercritical states only. Remarkably this procedure also predicts liquid-gas coexistence and associated phenomena.
@@ II.1 Overview of classical density functional theory
Classical DFT [8, 10, 11] ascertains the existence and uniqueness of the grand potential density functional, $\Omega([\rho],T,\mu)$, which consists of ideal, excess (over ideal gas), external, and chemical contributions according to the sum:
@@
where the position integrals run over the system volume $V$. Here and throughout we indicate functional relationships by square brackets. The ideal gas free energy functional is known exactly as $F_{\rm id}([\rho],T)=k_{B}T\int\mathrm{d}{\mathbf{r}}\,\rho({\mathbf{r}})[\ln(\rho({\mathbf{r}})\Lambda^{d})-1]$, where $\Lambda$ is the thermal de Broglie wavelength (which we will set to the particle size below). The excess free energy functional $F_{\rm exc}([\rho],T)$ accounts for the effects of the non-vanishing interparticle interactions. Crucially, instead of operating only on the true equilibrium density profile $\rho_{0}({\mathbf{r}})$, the grand potential density functional (1) accepts any general “test” function profile $\rho({\mathbf{r}})$ that does not need to have particular physical significance for the system at hand. Identifying the true equilibrium density profile $\rho_{0}({\mathbf{r}})$ is ensured by the formally exact minimization principle $\Omega([\rho_{0}],T,\mu)\leq\Omega([\rho],T,\mu)$ and hence
@@
Figure 1: Left panel: training data are acquired via grand canonical MC simulations of the truncated LJ fluid in inhomogeneous planar environments at randomized chemical potential $\mu$ and temperature $T$. We show the thermodynamic state points of all contributing simulations. The symbol size indicates the lateral system length $L$, which is also varied randomly. Small symbols refer to values of $L$ closer to $5\sigma$ and big symbols to values where $L$ is closer to $20\sigma$. The binodal and critical point taken from Ref. [57] are shown in gray. The critical temperature $k_{B}T_{c}/\varepsilon\approx 1.188$ provides an indicator of where we might ‘expect’ to find liquid-gas phase separation. Note that the systems are inhomogeneous, thus rendering the value of $\mu$ inconclusive for determining the emerging phases. Middle panel: an example from the training set. The randomly generated planar external potential $V_{\mathrm{ext}}(x)$ (black curve, hard walls indicated in gray) creates an inhomogeneous density profile $\rho(x)$ (blue), which is sampled in the simulation. The one-body direct correlation function $c_{1}(x)$ (orange) follows pointwise from Eq. (7). Right panel: a neural network is trained to extract and represent the underlying functional relationship $c_{1}(x;[\rho],T)$ in a local manner [25]. The parametric temperature dependence and the finite size scaling are taken into account via additional nodes in the input layer. We also indicate in the left panel the temperature cutoff of $k_{B}T/\varepsilon=1.3$ (horizontal dotted pink line) for the case of purely supercritical training, where only simulations with higher temperature contribute—see text.
@@ II.2 Generation of training data
Training data are acquired in grand canonical Monte Carlo (MC) simulation runs [9] with randomly generated inhomogeneous potential energy landscapes in planar geometry, i.e. $V_{\mathrm{ext}}(\mathbf{r})=V_{\mathrm{ext}}(x)$ [25]. We further randomize the thermodynamic state point as specified by the chemical potential $\mu$ and the temperature $T$, which are chosen uniformly within the ranges $-7<\mu/\varepsilon<4$ and $21<k_{B}T/\varepsilon<$ 2. For ease of sampling and histogram construction, the system length in the (inhomogeneous) $x$-direction is kept fixed at a value of $L_{x}=20\sigma$. To alleviate finite size effects, at least to some extent, we also vary the lateral system lengths $L_{y}=L_{z}=L$ uniformly within the interval $205<L/\sigma<$ 20. The advantages and limitations of this procedure, in particular regarding the resulting behaviour in the vicinity of the critical point, are described in Sec. III.5.
@@
Fig. 1 (left panel) shows the thermodynamic state points of all training simulations. Note that we also show the coexistence curve, in the $(\mu,T)$ plane, obtained in MC simulations by Wilding [57]. The critical temperature $k_{B}T_{c}/\varepsilon=1.188$ provides an important indication of where we might hope to find phase separation using our present neural functional. As an example of a simulation within the training set, namely at $\mu/\varepsilon\approx-0.17$ and $k_{B}T/\varepsilon\approx 1.76$, we show the relevant inhomogeneous one-body profiles $\rho(x)$ and $c_{1}(x)$ for a particular shape of $V_{\mathrm{ext}}(x)$. In the right panel of Fig. 1, we show a schematic illustration of the neural density functional mapping, which we describe in the following.
@@ II.3 Neural network and training procedures
To incorporate the parametric dependence of one-body direct correlation functions on temperature $T$ and lateral system size $L$, additional nodes in the input layer are provided, which accept the respective scalar values as indicated in Fig. 1. As a technical detail, we input $1/L$ instead of $L$, which allows us to set $01/L=$ 0 during inference to extrapolate to large lateral system sizes. We argue that the variability in $L$ has advantages over training with fixed lateral system size, although it may not account for the true finite size scaling behaviour due to insufficient information in the training data (see also Sec. III.5 for limitations in the critical region). Nevertheless, one may hope to avoid ingraining the specific finite size effects of a particular choice of $L$ via this procedure.
Investigating the extrapolation capabilities of the neural network is important for practical applications but also from a conceptual point of view. Recall that the neural functional framework relies upon extracting a functional mapping from reference data obtained for inhomogeneous equilibrium fluids, which satisfy the minimization principle (2) by definition. However, the underlying functional relationship might be much more general; it is not restricted a priori to true equilibrium density profiles, thereby raising profound mathematical questions about the existence of a unique continuation. In order to scrutinize this problem from a data-driven perspective, and in particular to show how much can be learned in the absence of any possible input information about coexistence, we train a second neural functional on the basis of supercritical data only by excluding simulations where $k_{B}T/\varepsilon<1.3$.
@@ II.4 Neural density functional theory and functional calculus
From a fundamental point of view, the availability of the entire density functional relationship $c_{1}(\mathbf{r};[\rho],T)$ suffices in principle to predict the full structural and thermodynamic behaviour of any fluid model. We lay out in this section common theoretical and numerical methods, which are taylored to the neural correlation functional and which differ in some aspects from the usual treatment of analytical (meaning explicit) free energy density functionals.
@@
which emerges from the minimization principle (2) and which determines the density profile $\rho(\mathbf{r})$ self-consistently. Eq. (8) can be solved efficiently, e.g. with a standard mixed Picard iteration, allowing for vast parameter studies, thereby commonly outperforming many-body simulation techniques by orders of magnitude in computational cost. The entirety of the nontrivial interparticle correlation effects are captured in the one-body direct correlation functional $c_{1}(\mathbf{r};[\rho],T)$, which are crucial in determining the resulting equilibrium state. Instead of approximating $c_{1}(\mathbf{r},[\rho],T)$ analytically, see e.g. the mean-field functional (4) that yields the one-body correlation functional explicitly upon functional differentiation (5), the machine learning routine in Secs. II.2 and II.3 provides a direct neural representation of this central functional mapping, which can be readily utilized in Eq. (8).
@@
can be evaluated efficiently from a computational representation of $c_{1}(\mathbf{r};[\rho],T)$ with reverse mode automatic differentiation (autodiff) [44]. This technique is particularly suited to our neural-network-based description of $c_{1}(\mathbf{r};[\rho],T)$, as autodiff is paramount to machine learning, specifically being the central mechanism for the backpropagation of errors during training [63]. As such, machine learning libraries come with ready-to-use implementations that make autodiff available as an efficient high-level operation, which we leverage for the evaluation of Eq. (9).
Contrary to standard analytical DFT approaches, which usually commence by expressing the excess free energy $F_{\mathrm{exc}}([\rho],T)$ as a density functional, e.g. in the form of the mean-field treatment (4), our starting point is given by the neural representation of the one-body direct correlation functional, which emerges formally as the functional derivative (5). Nevertheless, the free energy is pertinent both in its mathematical role as a generating functional as well as for the calculation of physical quantities such as the equation of state and the surface tension (see Sec. III.4). For evaluating $F_{\mathrm{exc}}([\rho],T)$ given a (neural) functional $c_{1}(\mathbf{r};[\rho],T)$, we utilize functional line integration [11, 25, 26, 27], which constitutes the formal inverse of the functional derivative (5). Making the functional line integral explicit via the linear parametrization of the density profile $\rho_{a}(\mathbf{r})=a\rho(\mathbf{r})$, $0\leq a\leq 1$, gives the expression [11, 25]

### v2

@@ Abstract
We use supervised machine learning together with the concepts of classical density functional theory to investigate the effects of interparticle attraction on the pair structure, thermodynamics, bulk liquid-gas coexistence, and associated interfacial phenomena in many-body systems. Local learning of the one-body direct correlation functional is based on Monte Carlo simulations of inhomogeneous systems with randomized thermodynamic conditions, randomized planar shapes of the external potential, and randomized box sizes. Focusing on the prototypical Lennard-Jones system, we test predictions of the resulting neural attractive density functional across a broad spectrum of physical behavior associated with liquid-gas phase coexistence in bulk and at interfaces. We analyse the bulk radial distribution function $g(r)$ obtained from automatic differentiation and the Ornstein-Zernike route and determine i) the Fisher-Widom line, i.e. the crossover of the asymptotic (large distance) decay of $g(r)$ from monotonic to oscillatory, ii) the (Widom) line of maximal correlation length, iii) the line of maximal isothermal compressibility and iv) the spinodal by calculating the poles of the structure factor in the complex plane. The bulk binodal and the density profile of the free liquid-gas interface are obtained from density functional minimization and the corresponding surface tension from functional line integration. We also show that the neural functional describes accurately the phenomena of drying at a hard wall and of capillary evaporation for a liquid confined in a slit pore. Our neural framework yields results that improve significantly upon standard mean-field treatments of interparticle attraction. Comparison with independent simulation results demonstrates a consistent picture of phase separation even when restricting the training to supercritical states only. We argue that phase coexistence and its associated signatures can be discovered as emerging phenomena via functional mappings and educated extrapolation.
@@ I Introduction
The emergence of two or more distinct thermodynamic stable phases on varying thermodynamic conditions is arguably one of the most striking phenomena in statistical mechanics, whether this occurs in bulk or at interfaces, in pure or in multi-component systems. Unsurprisingly, the recent surge in the use of machine-learning techniques in physics [1, 2] has focused on the prototypical (lattice-based) Ising model for developing appropriate techniques and strategies to investigate phase-separating systems. Such work includes the finite-size analysis in neural network classification of critical phenomena [3] and mapping out phase diagrams with generative classifiers [4]. Detecting the approach to a critical point also constitutes a central task in more general dynamical systems, which commonly requires the application of advanced computational methods [5, 6, 7].
@@
As DFT is an exact formulation of the many-body statistical mechanics, the choice of the excess free energy functional is the only approximation that enters a given study. Once it is specified consistent and complete investigation of a wide variety of properties can be made. Computational limitations, leaving aside some intricacies of implementing nonlocal treatments for hard spheres [13, 14], are minor, certainly for systems which exhibit planar or spherical symmetry. Hence wide parameter sweeps and close monitoring of the effects of small changes in control parameters, especially near phase coexistence, are readily carried out enabling the investigation of subtle phenomena such as phase transitions at substrates and in confinement.
@@
Machine learning provides a very different perspective for addressing the limitations of analytical approaches (meaning writing down explicit free energy functionals) by incorporating quasiexact simulation reference data for the construction of functional relationships, as was pursued for the classical [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34] and quantum (electronic) worlds [35, 36, 37, 38, 39, 40, 41, 42]. The review by Simon and Oettel [31] gives a valuable overview of very recent applications of machine-learning techniques within classical DFT and addresses methodological connections to electronic DFT and to nonequilibrium systems. Amongst the different approaches [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 35, 36, 37, 38, 39, 40, 41, 42, 43, 33, 34] that were put forward, the neural functional theory based on local one-body learning [24, 25, 26, 27, 28, 29, 33] in inhomogeneous training systems has proved to be a versatile and highly accurate tool, both in equilibrium [25, 26, 27, 28, 33] as well as for microscopically resolved nonequilibrium flow problems [24, 29]. In the latter case the required functional relationships are those of power functional theory [44, 45].
@@
The neural functional approach was investigated in great detail for models with purely repulsive (hard core) potentials. It provides an excellent approximation for Percus’ exact free energy functional for hard rods in one dimension [26] and was shown to constitute a clear improvement [25] over the already highly accurate White-Bear Mk. II version of fundamental-measure theory [56, 14] for hard spheres in three dimensions. Importantly, the neural functional method is not restricted to hard cores; it applies to general interatomic potentials which may also include attraction. That the neural functional can treat attraction successfully was illustrated in determinations of the structure of the Lennard-Jones fluid at a fixed supercritical temperature [25, 28]; this constitutes an important test case [23].
These recent investigations did not address the fundamental issue of how the presence of a phase transition might be accounted for within the framework of a neural density functional. Here we focus on the liquid-gas transition which is a basic manifestation of the presence of interparticle attraction and seek to assess whether the neural functional can describe i) phase coexistence and the approach to the associated critical point, ii) surface tension and density profiles of the liquid-gas interface, iii) drying and capillary evaporation transitions that occur at subcritical temperatures and iv) how accurately the approach performs for both bulk and interfacial properties. To the best of our knowledge, none of these issues were addressed in previous machine-learning investigations. Specifically, we extend the neural methodology [25, 26] by introducing thermal training and investigate whether neural functionals can describe physical phenomena that occur in the three-dimensional Lennard-Jones system in sub- as well as supercritical regions of the phase diagram. The answer is emphatically: yes.
@@
All subsequent results originate from this neural functional method and are described in Sec. III. An account of the emerging bulk pair correlation structure is given in Sec. III.1. Results for the lines in the phase diagram where the isothermal compressibility and the true correlation length are maximal, together with the Fisher-Widom line and the spinodal, are presented Sec. III.2. Our results for the liquid-gas binodal and estimate of the critical point are described in Sec. III.3. The bulk equation of state, liquid-gas density profiles, and the corresponding surface tension are laid out in Sec. III.4. A description of the divergence of the correlation length in the critical region and corresponding Ornstein-Zernike plots are given in Sec. III.5. Results for inhomogeneous fluids, that describe our predictions for drying at a hard wall, for capillary evaporation in a slit pore and for the corresponding behavior of locally resolved density fluctuations, are presented in Sec. III.6. In Sec. III.7 we compare with results obtained from a neural functional trained with data from supercritical states only. Remarkably this procedure also predicts liquid-gas coexistence and associated phenomena.
@@ II.1 Overview of classical density functional theory
Classical DFT [8, 10, 11] ascertains the existence and uniqueness of the grand potential density functional, $\Omega([\rho],T,\mu)$, which consists of ideal, excess (over ideal gas), external, and chemical contributions according to the sum
@@
where the position integrals run over the system volume $V$. Here and throughout we indicate functional relationships by square brackets. The ideal gas free energy functional is known exactly as $F_{\rm id}([\rho],T)=k_{B}T\int\mathrm{d}{\mathbf{r}}\,\rho({\mathbf{r}})[\ln(\rho({\mathbf{r}})\Lambda^{d})-1]$, where $\Lambda$ is the thermal de Broglie wavelength (which we will set to the particle size below). The excess free energy functional $F_{\rm exc}([\rho],T)$ accounts for the effects of the nonvanishing interparticle interactions. Crucially, instead of operating only on the true equilibrium density profile $\rho_{0}({\mathbf{r}})$, the grand potential density functional (1) accepts any general “test” function profile $\rho({\mathbf{r}})$ that does not need to have particular physical significance for the system at hand. Identifying the true equilibrium density profile $\rho_{0}({\mathbf{r}})$ is ensured by the formally exact minimization principle $\Omega([\rho_{0}],T,\mu)\leq\Omega([\rho],T,\mu)$ and hence
@@
Figure 1: Left panel: training data are acquired via grand canonical MC simulations of the truncated LJ fluid in inhomogeneous planar environments at randomized chemical potential $\mu$ and temperature $T$. We show the thermodynamic state points of all contributing simulations. The symbol size indicates the lateral system length $L$, which is also varied randomly. Small symbols refer to values of $L$ closer to $5\sigma$ and big symbols to values where $L$ is closer to $20\sigma$. The binodal and critical point taken from Ref. [59] are shown in gray. The critical temperature $k_{B}T_{c}/\varepsilon\approx 1.188$ provides an indicator of where we might ‘expect’ to find liquid-gas phase separation. Note that the systems are inhomogeneous, thus rendering the value of $\mu$ inconclusive for determining the emerging phases. Middle panel: an example from the training set. The randomly generated planar external potential $V_{\mathrm{ext}}(x)$ (black curve, hard walls indicated in gray) creates an inhomogeneous density profile $\rho(x)$ (blue), which is sampled in the simulation. The one-body direct correlation function $c_{1}(x)$ (orange) follows pointwise from Eq. (7). Right panel: a neural network is trained to extract and represent the underlying functional relationship $c_{1}(x;[\rho],T)$ in a local manner [25]. The parametric temperature dependence and the finite-size scaling are taken into account via additional nodes in the input layer. We also indicate in the left panel the temperature cutoff of $k_{B}T/\varepsilon=1.3$ (horizontal dotted pink line) for the case of purely supercritical training, where only simulations with higher temperature contribute—see text.
@@ II.2 Generation of training data
Training data are acquired in grand canonical Monte Carlo (MC) simulation runs [9] with randomly generated inhomogeneous potential energy landscapes in planar geometry, i.e. $V_{\mathrm{ext}}(\mathbf{r})=V_{\mathrm{ext}}(x)$ [25]. We further randomize the thermodynamic state point as specified by the chemical potential $\mu$ and the temperature $T$, which are chosen uniformly within the ranges $-7<\mu/\varepsilon<4$ and $21<k_{B}T/\varepsilon<$ 2. For ease of sampling and histogram construction, the system length in the (inhomogeneous) $x$-direction is kept fixed at a value of $L_{x}=20\sigma$. To alleviate finite-size effects, at least to some extent, we also vary the lateral system lengths $L_{y}=L_{z}=L$ uniformly within the interval $205<L/\sigma<$ 20. The advantages and limitations of this procedure, in particular regarding the resulting behavior in the vicinity of the critical point, are described in Sec. III.5.
@@
Fig. 1 (left panel) shows the thermodynamic state points of all training simulations. Note that we also show the coexistence curve (gray line) in the $(\mu,T)$ plane, obtained in grand canonical MC simulations by Wilding [59]. This ends at the critical temperature $k_{B}T_{c}/\varepsilon=1.188$ which provides an important indication of where we might hope to find phase separation using our present neural functional. As an example of a simulation within the training set, namely at $\mu/\varepsilon\approx-0.17$ and $k_{B}T/\varepsilon\approx 1.76$, we show the relevant inhomogeneous one-body profiles $\rho(x)$ and $c_{1}(x)$ for a particular shape of $V_{\mathrm{ext}}(x)$. In the right panel of Fig. 1, we show a schematic illustration of the neural density functional mapping, which we describe in the following.
@@ II.3 Neural network and training procedures
To incorporate the parametric dependence of one-body direct correlation functions on temperature $T$ and lateral system size $L$, additional nodes in the input layer are provided, which accept the respective scalar values as indicated in Fig. 1. As a technical detail, we input $1/L$ instead of $L$, which allows us to set $01/L=$ 0 during inference to extrapolate to large lateral system sizes. We argue that the variability in $L$ has advantages over training with fixed lateral system size, although it may not account for the true finite-size scaling behavior due to insufficient information in the training data (see also Sec. III.5 for limitations in the critical region). Nevertheless, one may hope to avoid ingraining the specific finite-size effects of a particular choice of $L$ via this procedure.
Investigating the extrapolation capabilities of the neural network is important for practical applications but also from a conceptual point of view. Recall that the neural functional framework relies upon extracting a functional mapping from reference data obtained for inhomogeneous equilibrium fluids, which satisfy the minimization principle (2) by definition. However, the underlying functional relationship might be much more general; it is not restricted a priori to true equilibrium density profiles, thereby raising profound mathematical questions about the existence of a unique continuation. In order to scrutinize this problem from a data-driven perspective, and in particular to show how much can be learned in the absence of any possible input information about coexistence, we train a second neural functional on the basis of supercritical data only by excluding simulations where $k_{B}T/\varepsilon<1.3$—see horizontal pink line in Fig. 1.
@@ II.4 Neural density functional theory and functional calculus
From a fundamental point of view, the availability of the entire density functional relationship $c_{1}(\mathbf{r};[\rho],T)$ suffices in principle to predict the full structural and thermodynamic behavior of any fluid model. We lay out in this section common theoretical and numerical methods, which are taylored to the neural correlation functional and which differ in some aspects from the usual treatment of analytical (meaning explicit) free energy functionals.
@@
which emerges from the minimization principle (2) and which determines the density profile $\rho(\mathbf{r})$ self-consistently. Eq. (8) can be solved efficiently, e.g. with a standard mixed Picard iteration, allowing for vast parameter studies, thereby commonly outperforming many-body simulation techniques by orders of magnitude in computational cost. The entirety of the nontrivial interparticle correlation effects are captured in the one-body direct correlation functional $c_{1}(\mathbf{r};[\rho],T)$, which are crucial in determining the resulting equilibrium state. Instead of approximating $c_{1}(\mathbf{r};[\rho],T)$ analytically, see e.g. the mean-field functional (4) that yields the one-body direct correlation functional explicitly upon functional differentiation (5), the machine-learning routine in Secs. II.2 and II.3 provides an immediate neural representation of this central functional mapping, which can be readily utilized in Eq. (8).
@@
can be evaluated efficiently from a computational representation of $c_{1}(\mathbf{r};[\rho],T)$ with reverse mode automatic differentiation (autodiff) [46]. This technique is particularly suited to our neural-network-based description of $c_{1}(\mathbf{r};[\rho],T)$, as autodiff is paramount to machine learning, specifically being the central mechanism for the backpropagation of errors during training [65]. As such, machine-learning libraries come with ready-to-use implementations that make autodiff available as an efficient high-level operation, which we leverage for the evaluation of Eq. (9).
Contrary to standard analytical DFT approaches, which usually commence by expressing the excess free energy $F_{\mathrm{exc}}([\rho],T)$ as an explicit density functional, e.g. in the form of the mean-field treatment (4), our starting point is given by the neural representation of the one-body direct correlation functional, which emerges formally as the functional derivative (5). Nevertheless, the free energy is pertinent both in its mathematical role as a generating functional as well as for the calculation of physical quantities such as the equation of state and the surface tension (see Sec. III.4). For evaluating $F_{\mathrm{exc}}([\rho],T)$ given a (neural) functional $c_{1}(\mathbf{r};[\rho],T)$, we utilize functional line integration [11, 25, 26, 27], which constitutes formally integrating the functional derivative (5). Making the functional line integral explicit via the linear parametrization of the density profile $\rho_{a}(\mathbf{r})=a\rho(\mathbf{r})$, $0\leq a\leq 1$, gives the expression [11, 25]